The acceleration of a particle at any time is given by a = 12e

3t

i − 8sin2tj +


4tk. If the velocity is zero at t = 0, find velocity.


1
Expert's answer
2022-04-18T01:45:39-0400

The acceleration of a particle at any time is given by

"\bold{a}=12e^{3t}\bold{i}-8\sin2t\bold{j} +4t\bold{k}" .

Velocity:

"\bold v=\int \bold a dt=\int (12e^{3t}\bold{i}-8\sin2t\bold{j} +4t\bold{k})dt=\\
\frac{12}{3}e^{3t}\bold{i}+\frac 82\cos2t\bold{j} +\frac 42 t^2\bold{k}+C_i\bold{i}+C_j\bold{j} +C_k\bold{k}=\\
4e^{3t}\bold{i}+4\cos2t\bold{j} +2 t^2\bold{k}+C_i\bold{i}+C_j\bold{j} +C_k\bold{k}",

where "C_i,C_j,C_k" are const.

"\bold v(0)=4\bold{i}+4\bold{j} +0\bold{k}+C_i\bold{i}+C_j\bold{j} +C_k\bold{k}=\bold 0"

"C_i=-4" ; "C_j=-4" ; "C_k=0" .

Answer:

"\bold v=4(e^{3t}-1)\bold{i}+4(\cos2t-1)\bold{j} +2 t^2\bold{k}"